# Leveraging Discrete Choice Experiments for User-Centric Requirements Prioritization in mHealth Applications

### v1

@@ 5.1 Implications for mHealth Developers
Our DCE study explored participants’ preferences for various attributes in mHealth app adaptation design and their willingness to make trade-offs, focusing on different subgroups such as gender, coping strategies, and health conditions. Figure 4 visually illustrates the trade-offs among different attributes and delineates how these trade-offs vary across different subgroups. The findings highlight app usability as the most critical attribute across all subgroups, underscoring its pivotal role in users’ decision-making processes regarding adaptation design. However, the study also reveals that providing users with controllability over adaptations, such as the ability to remove specific features, revert changes, or adjust adaptation levels, can mitigate potential usability trade-offs. User-controlled adaptation fosters a sense of ownership and agency, leading to increased engagement and motivation to utilize the system effectively [sundar2010personalization, sundar2008self]. This suggests that while usability remains paramount, empowering users with control over adaptations can enhance acceptance and satisfaction, even in the presence of usability challenges.
@@
Incorporating caregivers into mHealth systems can influence user preferences for app adaptations, as patients can delegate certain responsibilities to caregivers. However, involving caregivers can also create a negative attitude toward adaptations. For example, apps like Health2Sync [^7] https://www.health2sync.com/ and LibreLinkUp [^8] https://www.librelinkup.com/ include caregivers in chronic disease management but limit their role to monitoring, without allowing them to control or customize app features for patients. This suggests that while caregiver involvement can be beneficial, practitioners should carefully design adaptation mechanisms to maximize the benefits. In particular, permit caregivers to aid in distinct tasks (such as personalizing reminders or monitoring health information) instead of granting patients sole control over the application’s capabilities. Ensure that adaptations do not degrade the app’s usability, as patients are unwilling to sacrifice ease of use for caregiver involvement.
@@
The frequency and criticality of app functionalities play a significant role in determining user preferences for adaptations. Users are less inclined to accept adaptations for frequently used or essential functionalities, as these tasks require consistency and reliability. This aligns with findings from lavie2010benefits, who demonstrated that the proportion of routine versus non-routine tasks influences the effectiveness of AUIs. Specifically, adaptations are less desirable for tasks that are performed regularly or are critical to the user’s workflow. For mHealth applications, where users often engage in a mix of routine and non-routine tasks [bunt2004role], this insight is particularly relevant. Practitioners should avoid introducing lots of adaptations for frequently used or essential functionalities to maintain user trust and efficiency. Focus on adapting features that are used infrequently or are less critical, where users may be more open to changes. Ensure that adaptations do not disrupt the usability of core functionalities, as these are key to user satisfaction and engagement.
@@
Patients prefer adaptations that occur infrequently and only when necessary. Once an adaptation aligns with their capabilities and needs, they see little value in frequent changes [wang2024adaptiveicse]. This suggests that adaptations should focus on essential functionalities and aim to enhance usability without introducing unnecessary disruptions. Importantly, users are willing to tolerate lower usability if it means avoiding frequent adaptations, particularly for less-used features. To create effective adaptive mHealth applications, practitioners should minimize adaptation frequency by designing changes to occur only when absolutely necessary. Practitioners should also focus on core functionalities by prioritizing adaptations for essential tasks that significantly improve usability or accessibility. Additionally, it is crucial to avoid over-adaptation by resisting frequent changes. Frequent adaptations can lead to frustration and reduced trust in the app.
@@
Patients prefer adaptations that are small-scale and incremental, ensuring the app remains familiar even after adaptations are implemented. For example, they are more comfortable with adaptations that modify specific components, such as the navigation bar or the order of data entry fields, rather than large-scale overhauls of the interface. Notably, patients are willing to tolerate slightly lower usability if it means preserving familiarity and avoiding drastic changes. Practitioners should make small, specific adaptations to app components rather than extensive modifications. Maintain the app’s structure and flow to keep it familiar to users, recognizing their preference for familiarity and consistency, even at the expense of some usability improvements.
@@
Summary. When designing adaptive mHealth apps without caregiver involvement, practitioners should focus on less frequently used functionalities for adaptations, as these are less likely to disrupt the user experience. For example, adaptations could target features like data export settings or advanced analytics, which users interact with infrequently. These adaptations should provide essential assistance while being infrequent and small-scale to avoid overwhelming the user. When caregivers are involved in mHealth app usage, the design should prioritize collaborative functionality while maintaining a balance between patient autonomy and caregiver support. For example, caregivers could be given access to monitor health data or set reminders, but patients should retain control over critical app customizations to ensure usability and comfort. For the rest of the factors for adaptation design. The design should follow the recommendations outlined earlier, such as focusing on small-scale, infrequent changes and prioritizing less frequently used functionalities. To address the varying frequency of functionality usage, a practical design approach is to introduce adaptations for less frequently used features during the first login. This allows users to customize these features upfront, ensuring personalization without the need for frequent changes later. For frequently used functionalities, such as daily health tracking or medication reminders, it is better to let users interact with the app for a period of time before offering adaptations.
@@ 5.1.1 Heterogeneous preferences across different user groups
Gender. Our analysis indicates that while men prioritize usability, women place greater emphasis on adaptability and customization, underscoring the importance of gender-specific design considerations (See Figure 4). According to Hofstede’s Masculinity vs. Femininity dimension, masculine cultures tend to favor practical problem-solving strategies, whereas feminine cultures place greater value on emotional support and personalized care [ahmed2024cultural]. These preferences align with the behavioral differences observed in the usage of mobile applications. For example, women tend to use smartphones for longer periods than men [andone2016age], and their app-related gratifications are more focused on enjoyment and goal-setting for physical activity tracking applications. In contrast, men are more likely to engage in live tracking and sharing of their results [klenk2017gender]. To enhance usability for male users, designs should prioritize streamlined interfaces, performance-oriented analytics, and quick-access shortcuts, ensuring efficiency without excessive complexity. In contrast, female users benefit from greater adaptability through personalized settings, context-aware assistance, motivational feedback, and progress-tracking tools that align with their preference for social and emotional engagement. Meanwhile, men place less value on caregiver involvement, so the design of mHealth applications should focus on enhancing user autonomy by providing intuitive self-management tools, clear guidance for independent use, and seamless access to necessary resources without relying on external support.
@@
Age. A clear pattern emerged among older individuals, who were more likely to trade off usability for less frequent and small-scale adaptations while placing less emphasis on controllability compared to younger users (See Figure 4). This preference may be due to age-related memory decline and limited familiarity with technology, which can increase susceptibility to confusion and cognitive overload [isakovic2016usability, wilson2021barriers, li2021design]. Frequent changes or an abundance of options within an application can feel overwhelming for older users, making stability and simplicity in design critical [garcia2018enabling]. The current mHealth landscape is deficient in user-friendly tools and services, with elderly individuals particularly facing usability challenges when using mHealth technology [wildenbos2019mobile]. Extensive research has been conducted on designing mHealth applications to better suit elderly users, focusing on features such as simple and consistent navigation [wildenbos2015framework, kalimullah2017influence], presenting all essential content on one screen and using a simple and familiar language [wildenbos2019mobile, pan2021perception]. Additionally, cultural considerations in UI design have been highlighted as critical for improving usability for older adults [alsswey2020elderly]. When adapting mHealth applications for elderly users, it is recommended that changes occur infrequently and only modify specific parts of the interface, ensuring that the overall app remains familiar and easy to navigate.
@@
Coping mechanism. Patients must continually develop strategies to maintain their physical, emotional, and spiritual health due to the long-lasting nature of chronic disease [bussing2010adaptive], making coping with the disease an ongoing process. Some individuals cope by ignoring distressing situations through avoidance and denial, while others use positive strategies to alleviate stress and restore normal functioning [folkman1980analysis]. Coping strategies affect consumers’ beliefs about new technology, which subsequently mediate the impact of these strategies on their attitude towards adoption [cui2009consumers]. In this research, the coping strategies shaped the preferences of the participants regarding the adaptation design. For individuals with positive coping styles, the design should enable collaboration by incorporating caregiver involvement, such as shared access to health data and communication tools, while also offering customization settings to provide greater control over app functionality. Conversely, for those with negative coping styles, the focus should be on simplicity and stability, with infrequent changes and small-scale adaptations to maintain familiarity and minimize disruptions (See Figure 4). Caregiver integration should be optional, respecting their preference for independent usage, while subtle prompts and non-intrusive notifications can gently encourage engagement.
@@
Health conditions. Chronic diseases impose diverse physical, psychological, and mental effects on individuals [harvey2012future, di2019chronic]. Research has shown that changes in health conditions over time for those with chronic diseases significantly affect their use of technology, a crucial subject that remains underexplored [paymal2024good]. Health conditions played a critical role in shaping attribute preferences. Participants in good health prioritize controllability, reflecting a preference for autonomy, while those in poor health emphasize small-scale adaptations and caregiver involvement, indicating a need for support and precision (See Figure 4). The design of mHealth technology frequently does not account for the fluctuation of mHealth conditions [paymal2024good, mack2022chronically], and most research tends to focus on a single chronic condition, despite the increasing prevalence of multimorbidity over time [islam2014multimorbidity, sockolow2021integrative]. To accommodate participants with diverse health conditions, it is essential to tailor adaptations. For healthy users, emphasizing controllability is crucial through customization settings, manual override functionalities, and interfaces that allow users to fine-tune their experiences. and for some frequent In contrast, designs such as the simplicity of design modifications on a small scale and straightforward caregiver access for supervision and assistance offer advantages to those with poor health. Since health conditions change over time, mHealth applications should include dynamic adaptation mechanisms that modify features according to current health conditions.
@@ 5.2.1 Using DCEs to Design mHealth Applications
mHealth app growth has surpassed evidence of their effectiveness, creating a gap in understanding the optimal design to meet user needs [kumar2013mobile]. Many mHealth interventions are designed around existing healthcare systems, often without sufficient input from end users [mccurdie2012mhealth]. This user-centered design shortfall can lead to apps that are hard to use or do not meet users’ specific needs, prove ineffective, or, in some cases, lead to adverse outcomes [brown2013assessment, schnall2016user]. In a healthcare system already burdened by suboptimal outcomes and rising costs, premature adoption of untested mHealth technologies risks exacerbating these challenges. Using a DCE can inform the design of mHealth applications by identifying key features that promote user uptake and engagement [jonker2020covid, simblett2023patient]. Unlike traditional methods, the experimental nature of DCEs enables researchers to systematically vary attributes/features under controlled conditions, providing insight into the marginal effects of attribute/features changes on user choices [johnson2013constructing].
@@
By identifying high-priority features early in the design process, DCEs can help reduce development costs and ensure that resources are allocated to features that maximize user engagement and satisfaction. For example, in our DCE study, we found that involving caregivers in the app design significantly influenced patients’ perceptions of adaptations, highlighting the importance of considering interdependence between adaptation features and user roles. DCEs also provide a structured approach to understanding how different features interact and influence user preferences. For instance, in our DCE study, the level of controllability over the adaptations can have cascading effects on the app usability. By quantifying these trade-offs, DCEs enable researchers to make informed design decisions that balance competing priorities and align with user needs.
@@
However, while DCEs offer significant advantages, they also present challenges, such as the complexity of designing choice tasks and the need for large sample sizes. There are existing tools (e.g., Qualtric, Ngene [^9] https://www.choice-metrics.com/index.html, SPSS [^10] https://www.ibm.com/products/spss-statistics) and papers (e.g., [weber2021step], [bridges2011conjoint], [szinay2021understanding]) to teach how to do the DCE step by step. The methodology remains challenging, especially for those without prior experience, as the intricate process of creating realistic choice scenarios, ensuring attribute balance, and analyzing complex data poses significant barriers to entry for novice researchers. Future research could focus on developing tools to simplify the implementation of DCEs for researchers and practitioners, making the methodology more accessible to those without extensive experience. Exploring mixed methods that integrate DCEs with qualitative approaches, such as interviews or focus groups, could offer deeper insights into user preferences by allowing researchers to quantify trade-offs and comprehend the contextual factors and personal experiences influencing decisions.
@@ 5.2.2 Identifying and Meeting Non-Functional Requirements (NFRs)
Non-functional requirements (NFRs) are global constraints such as performance, security, and availability that software must meet, often across multiple system components [rosa2002processnfl]. These requirements play a critical role in ensuring software quality, user satisfaction, and project success. However, mismanagement of NFRs has been a major cause of many software project failures [ryan2000approach], often due to their inherent complexity, interdependencies, and conflicts. Various methods, such as fuzzy logic [saadatmand2015fuzzy], ontology-based techniques [mairiza2013conflict, shah2019ontological], and machine learning or deep learning [alashqar2022studying], automate the extraction, classification, and prioritization of NFRs from requirements documents. However, these approaches often fail to capture end-user preferences and priorities, limiting their effectiveness in designing user-centered systems. The question of how users perceive and manage these competing NFRs remains largely unexplored. For example, in mHealth applications, users may value privacy highly, but may be unwilling to sacrifice usability or performance for improved privacy [hamidi2018should]. DCEs offer a promising framework to address this gap by systematically quantifying user preferences for such trade-offs. By simulating real-world decision-making scenarios, DCEs allow researchers to uncover hidden priorities and inform design decisions that align with user needs. In addition, DCEs could address evolving challenges in NFR management, such as sustainability and ethical considerations, ensuring that software systems are not only functional, but also socially responsible.

### v2

@@ 5.1 Implications for mHealth Developers
Our DCE study explored participants’ preferences for various attributes in mHealth app adaptation design and their willingness to make trade-offs, focusing on different subgroups such as gender, coping strategies, and health conditions. Figure 4 visually illustrates the trade-offs between different attributes and delineates how these trade-offs vary between different subgroups. The findings highlight app usability as the most critical attribute across all subgroups, underscoring its pivotal role in users’ decision-making processes regarding adaptation design. However, the study also reveals that providing users with controllability over adaptations, such as the ability to remove specific features, revert changes, or adjust adaptation levels, can mitigate potential usability trade-offs. User-controlled adaptation fosters a sense of ownership and agency, leading to increased engagement and motivation to utilize the system effectively [87, 88]. While caregiver involvement can offer benefits, the observed negative attitudes indicate that adaptation mechanisms should be carefully designed to balance control and support.
@@
Incorporating caregivers into mHealth systems can influence user preferences for app adaptations, as patients can delegate certain responsibilities to caregivers. However, involving caregivers can also create a negative attitude toward adaptations. For example, apps like Health2Sync [^7] https://www.health2sync.com/ and LibreLinkUp [^8] https://www.librelinkup.com/ include caregivers in chronic disease management but limit their role to monitoring, without allowing them to control or customize app features for patients. This suggests that while caregiver participation can be beneficial, practitioners should carefully design adaptation mechanisms to maximize benefits. In particular, permit caregivers to aid in distinct tasks (such as personalizing reminders or monitoring health information) instead of granting patients sole control over the application’s capabilities. Ensure that adaptations do not degrade the app’s usability, as patients are unwilling to sacrifice ease of use for caregiver involvement.
@@
The frequency and criticality of the app functionalities play a significant role in determining user preferences for adaptations. Users are less inclined to accept adaptations for frequently used or essential functionalities, as these tasks require consistency and reliability. This aligns with the findings of Lavie and Meyer [60], who demonstrated that the proportion of routine versus non-routine tasks influences the effectiveness of AUIs. Specifically, adaptations are less desirable for tasks that are performed regularly or are critical to the user’s workflow. For mHealth applications, where users often engage in a mix of routine and non-routine tasks [18], this insight is particularly relevant. Practitioners should avoid introducing many adaptations for commonly used or essential functionalities to maintain user trust and efficiency. Focus on adapting features that are used infrequently or are less critical, where users may be more open to changes. Ensure that adaptations do not disrupt the usability of core functionalities, as these are key to user satisfaction and engagement.
@@
Patients prefer adaptations that occur infrequently and only when necessary. Once an adaptation aligns with their capabilities and needs, they see little value in frequent changes [98]. This suggests that adaptations should focus on essential functionalities and aim to enhance usability without introducing unnecessary disruptions. Importantly, users are willing to tolerate lower usability if it means avoiding frequent adaptations, particularly for less-used features. To create effective adaptive mHealth applications, practitioners should minimize adaptation frequency by designing changes to occur only when absolutely necessary. Practitioners should also focus on core functionalities by prioritizing adaptations for essential tasks that significantly improve usability or accessibility. Additionally, it is crucial to avoid over-adaptation by resisting frequent changes. Frequent adaptations can lead to frustration and reduced trust in the app.
@@
Patients prefer small-scale and incremental adaptations, ensuring that the app remains familiar even after adaptations are implemented. For example, they are more comfortable with adaptations that modify specific components, such as the navigation bar or the order of data entry fields, rather than large-scale overhauls of the interface. This finding suggests that familiarity may outweigh usability improvements for some users. Practitioners should make small, specific adaptations to app components rather than extensive modifications. Maintain the app’s structure and flow to keep it familiar to users, recognizing their preference for familiarity and consistency, even at the expense of some usability improvements.
@@
Summary. When designing adaptive mHealth apps without caregiver involvement, practitioners should focus on less frequently used functionalities for adaptations, as these are less likely to disrupt the user experience. For example, adaptations could target features like data export settings or advanced analytics, which users interact with infrequently. These adaptations should provide essential assistance while being infrequent and small-scale to avoid overwhelming the user. When caregivers are involved in mHealth app usage, the design should prioritize collaborative functionality while maintaining a balance between patient autonomy and caregiver support. For example, caregivers could have access to monitor health data or set reminders, but patients should retain control over critical app customizations to ensure usability and comfort. For the rest of the factors for adaptation design. The design should follow the recommendations outlined earlier, such as focusing on small-scale, infrequent changes and prioritizing less frequently used functionalities. To address the varying frequency of functionality usage, a practical design approach is to introduce adaptations for less frequently used features during the first login. This allows users to customize these features upfront, ensuring personalization without the need for frequent changes later. For frequently used functionalities, such as daily health tracking or medication reminders, it is better to let users interact with the app for a period of time before offering adaptations.
@@ 5.1.1 Heterogeneous preferences across different user groups
Gender. Our analysis indicates that while men prioritize usability, women place greater emphasis on adaptability and customization, underscoring the importance of gender-specific design considerations (See Figure 4). According to Hofstede’s Masculinity vs. Femininity dimension, masculine cultures tend to favor practical problem-solving strategies, whereas feminine cultures place greater value on emotional support and personalized care [3]. These preferences align with the behavioral differences observed in the usage of mobile applications. For example, women tend to use smartphones for longer periods than men [10], and their app-related gratifications are more focused on enjoyment and setting goals for physical activity tracking applications. In contrast, men are more likely to engage in live tracking and sharing of their results [56]. To improve usability for male users, designs should prioritize streamlined interfaces, performance-oriented analytics, and quick-access shortcuts, ensuring efficiency without excessive complexity. In contrast, female users benefit from greater adaptability through personalized settings, context-aware assistance, motivational feedback, and progress-tracking tools that align with their preference for social and emotional engagement. Meanwhile, men place less value on caregiver involvement, so the design of mHealth applications should focus on improving user autonomy by providing intuitive self-management tools, clear guidance for independent use, and seamless access to necessary resources without relying on external support.
@@
Age. A clear pattern emerged among older individuals, who were more likely to trade off usability for less frequent and small-scale adaptations while placing less emphasis on controllability compared to younger users (See Figure 4). This preference may be due to age-related memory decline and limited familiarity with technology, which can increase susceptibility to confusion and cognitive overload [45, 103, 61]. Frequent changes or an abundance of options within an application can feel overwhelming for older users, making stability and simplicity in design critical [33]. The current mHealth landscape is deficient in user-friendly tools and services, with elderly people particularly facing usability challenges when using mHealth technology [101]. Extensive research has been conducted on designing mHealth applications to better suit elderly users, focusing on features such as simple and consistent navigation [102, 55], presenting all essential content on one screen and using a simple and familiar language [101, 70]. Additionally, cultural considerations in UI design have been highlighted as critical for improving usability for older adults [8]. Based on the observed preferences, mHealth applications for elderly users may benefit from infrequent, small-scale adaptations that modify only specific parts of the interface while preserving familiarity and ease of navigation.
@@
Coping mechanism. Patients must continually develop strategies to maintain their physical, emotional, and spiritual health due to the long-lasting nature of chronic disease [20], making coping with the disease an ongoing process. Some individuals cope by ignoring distressing situations through avoidance and denial, while others use positive strategies to alleviate stress and restore normal functioning [28]. Coping strategies affect consumers’ beliefs about new technology, which subsequently mediate the impact of these strategies on their attitude towards adoption [23]. Our findings suggest that coping strategies shape users’ preferences for adaptation design. For individuals with positive coping styles, adaptive mHealth designs can benefit from supporting collaboration through caregiver involvement, such as shared access to health data and communication tools, while also providing customization options that improve user control over app functionality. In contrast, the observed preferences of individuals with negative coping styles indicate a stronger need for simplicity and stability, favoring infrequent and small-scale adaptations that preserve familiarity and minimize disruption (see Figure 4). These patterns suggest that caregiver integration should remain optional, respecting preferences for independent use, while subtle prompts and non-intrusive notifications may help encourage engagement without imposing additional burden.
@@
Health conditions. Chronic diseases impose diverse physical, psychological, and mental effects on individuals [42, 26]. Research has shown that changes in health conditions over time for those with chronic diseases significantly affect their use of technology, a crucial subject that remains underexplored [71]. Health conditions played a critical role in shaping attribute preferences. Participants in good health prioritize controllability, reflecting a preference for autonomy, while those in poor health emphasize small-scale adaptations and caregiver involvement, indicating a need for support and precision (See Figure 4). The design of mHealth technology frequently does not account for the fluctuation of mHealth conditions [71, 62], and most research tends to focus on a single chronic condition, despite the increasing prevalence of multimorbidity over time [46, 85]. These findings suggest that adaptations should be tailored to accommodate users with diverse health conditions. For individuals in relatively good health, emphasizing controllability—through customization settings, manual override functionalities, and interfaces that allow users to fine‑tune their experiences—may better support autonomy. In contrast, the observed preferences of participants in poorer health indicate greater benefit from simple, small‑scale design modifications and straightforward caregiver access for supervision and assistance. As health conditions can change over time, these results further suggest that mHealth applications should incorporate dynamic adaptation mechanisms that adjust features in response to users’ current health status.
@@ 5.2.1 Using DCEs to Design mHealth Applications
The growth of mHealth apps has exceeded the evidence of their effectiveness, creating a gap in understanding the optimal design to meet user needs [58]. Many mHealth interventions are designed around existing healthcare systems, often without sufficient input from end users [65]. This user-centered design shortfall can lead to apps that are hard to use or do not meet users’ specific needs, prove ineffective, or, in some cases, lead to adverse outcomes [17, 82]. In a healthcare system already burdened by suboptimal outcomes and rising costs, premature adoption of untested mHealth technologies risks exacerbating these challenges. Using a DCE can inform the design of mHealth applications by identifying key features that promote user uptake and engagement [54, 84]. Unlike traditional methods, the experimental nature of DCEs enables researchers to systematically vary attributes under controlled conditions, providing insight into the marginal effects of attribute changes on user choices [50].
@@
By identifying high-priority features early in the design process, DCEs can help reduce development costs and ensure that resources are allocated to features that maximize user engagement and satisfaction. For example, in our DCE study, we found that involving caregivers in the app design significantly influenced patients’ perceptions of adaptations, highlighting the importance of considering the interdependence between adaptation features and user roles. DCEs also provide a structured approach to understanding how different features interact and influence user preferences. For instance, in our DCE study, the level of controllability over the adaptations can have cascading effects on the app usability. By quantifying these trade-offs, DCEs enable researchers to make informed design decisions that balance competing priorities and align with user needs.
@@
However, while DCEs offer significant advantages, they also present challenges, such as the complexity of designing choice tasks and the need for large sample sizes. There are existing tools (e.g., Qualtric, Ngene [^9] https://www.choice-metrics.com/index.html, SPSS [^10] https://www.ibm.com/products/spss-statistics) and papers (e.g., [99], [16], [89]) to teach how to do the DCE step by step. The methodology remains challenging, especially for those without prior experience, as the intricate process of creating realistic choice scenarios, ensuring attribute balance, and analyzing complex data poses significant barriers to entry for novice researchers. Future research could focus on developing tools to simplify the implementation of DCEs for researchers and practitioners, making the methodology more accessible to those without extensive experience. Exploring mixed methods that integrate DCEs with qualitative approaches, such as interviews or focus groups, could offer deeper insights into user preferences by allowing researchers to quantify trade-offs and comprehend the contextual factors and personal experiences influencing decisions.
@@ 5.2.2 Identifying and Meeting Non-Functional Requirements (NFRs)
Non-functional requirements (NFRs) are global constraints such as performance, security, and availability that software must meet, often across multiple system components [77]. These requirements play a critical role in ensuring software quality, user satisfaction, and project success. However, mismanagement of NFRs has been a major cause of many software project failures [78], often due to their inherent complexity, interdependencies, and conflicts. Various methods, such as fuzzy logic [81], ontology-based techniques [63, 83], and machine learning or deep learning [7], automate the extraction, classification, and prioritization of NFRs from requirements documents. However, these approaches often fail to capture end-user preferences and priorities, limiting their effectiveness in designing user-centered systems. The question of how users perceive and manage these competing NFRs remains largely unexplored. For example, in mHealth applications, users may value privacy highly, but may be unwilling to sacrifice usability or performance for improved privacy [39]. DCEs offer a promising framework to address this gap by systematically quantifying user preferences for such trade-offs. By simulating real-world decision-making scenarios, DCEs allow researchers to uncover hidden priorities and inform design decisions that align with user needs. In addition, DCEs could address evolving challenges in NFR management, such as sustainability and ethical considerations, ensuring that software systems are not only functional, but also socially responsible.